OpenStudy (anonymous):

Bill wants to factor x2 + 8x + 16 by grouping; however, Sara says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (anonymous):

@surry99

OpenStudy (surry99):

looks like you have some help on the way

OpenStudy (anonymous):

yes hopefully but can you stick around because i have two more? @surry99

OpenStudy (anonymous):

@surry99 im gonna post this one next can you review it for me please Given the function f(x) = 4(x + 1)2 − 3, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x−2) f(x) − 2 f(2x) 2•f(x)

OpenStudy (anonymous):

f(x) =4(x+1)^2-3

OpenStudy (anonymous):

where those question marks are that is a minus

OpenStudy (calculusfunctions):

"By grouping" implies the method of decomposition. Example: Factor\[x ^{2}+6x+9\]First we find that the two integers with a product of 9 and a sum of 6 are 3 and 3. Hence\[x ^{2}+6x +9\]\[=x ^{2}+3x +3x +9\]\[=x(x +3)+3(x +3)\]\[=(x +3)(x +3)\]\[=(x +3)^{2}\]

OpenStudy (calculusfunctions):

The other method is recognizing a trinomial as a perfect square trinomial. If given\[ax ^{2}+bx +c\]Where a and c a re both perfect squares and b (regardless of the sign) is twice the product of the square roots of a and c, then\[ax ^{2}+bx +c =(a +b)^{2}\]OR\[ax ^{2}-bx +c =(a -b)^{2}\]

OpenStudy (anonymous):

so the first one is an answer?

OpenStudy (calculusfunctions):

What do you mean by "the first one is an answer"?

OpenStudy (anonymous):

the grouping way?

OpenStudy (anonymous):

thats how i would do it

OpenStudy (calculusfunctions):

That is the grouping way but that was an example, not the specific answer to your question. I expect you to follow the example and solve your own questions.

OpenStudy (calculusfunctions):

Yes! That is how you would do it!

OpenStudy (anonymous):

ok can you look at the second one i posted please ima open a new question

OpenStudy (calculusfunctions):

Not able to read the entire question. I think you need to type it out in its entirety or provide a link if one exists.

OpenStudy (anonymous):

ok ima repost it ok!